OpenStudy (anonymous):

Find the arc length of the curve r=θ^3, 0≤θ≤π. Where are the tangent lines vertical? Horizontal?

OpenStudy (anonymous):

you need y = f(θ) and x = f(θ), do you know how to find them?

OpenStudy (anonymous):

I belive so. wouldn't the y function be cos(θ) - θsin(θ) and the x function be sinθ + θcos(θ)?

OpenStudy (anonymous):

Well, let's do the arc length first. Do you know the formula?

OpenStudy (anonymous):

\[\sqrt{dx + dy}\] ?

OpenStudy (anonymous):

dx and dy are squared though.

OpenStudy (anonymous):

well you could do it that way, but there is a better alternative. L = ∫sqrt(r^2 + (dr/dθ)&2) dθ. Look familiar?

OpenStudy (anonymous):

Not at all. Could you elaborate on that?

OpenStudy (anonymous):

typo.... L = ∫sqrt(r^2 + (dr/dθ)^2) dθ

OpenStudy (anonymous):

Ok, so it should look like this: \[\sqrt{r ^{2} + dr/dTheta^2}\] albeit with the integration, correct?

OpenStudy (anonymous):

yes. r = θ^3 so dr/dθ = 3θ^2 yes?

OpenStudy (anonymous):

Ok, that makes sense. So, can you explain why I need to find the arc length?

OpenStudy (anonymous):

well the question asks for arclength doesn't it?

OpenStudy (anonymous):

Sorry, you're correct. (Long day with a decent amount of medicine makes one a bit oblivious)

OpenStudy (anonymous):

so, when plug those in the formula what do you get?

OpenStudy (anonymous):

I ended up getting sqrt(theta^5 + 9theta^4). Right?

OpenStudy (anonymous):

uhm, what is (θ^3)^2 ?

OpenStudy (anonymous):

Ohh, I see what I did. That should be to the sixth.

OpenStudy (anonymous):

simplify further what do you get?

OpenStudy (anonymous):

theta to the third + 3(theta)^2

OpenStudy (anonymous):

O.O holy crap man. You can't do that!

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

sqrt(θ^6 + 9θ^4) = sqrt[θ^4 (θ^2 + 9)] <--- do you see what I did?

OpenStudy (anonymous):

Ok, I see what you did.

OpenStudy (anonymous):

ok, what did i do? and how do you simplify it further

OpenStudy (anonymous):

took a theta^4 out of both terms, and the quantity can be broken down into theta + 3 and theta -3

OpenStudy (anonymous):

Nooooooooooooooooooooo!!!!!! D: You told me the same mistake again \[\sqrt{a+b} \neq \sqrt{a} + \sqrt{b}\]

OpenStudy (anonymous):

*facepalm*

OpenStudy (anonymous):

hower, \[\sqrt{ab} = (\sqrt{a}) (\sqrt{b})\]

OpenStudy (anonymous):

*however*,...

OpenStudy (anonymous):

Using that formula above, can you simplify it further?

OpenStudy (anonymous):

Take the theta^4 and the quantity, separate them into two different square root functions, and go from there?

OpenStudy (anonymous):

yes, specifically simplified to what?

OpenStudy (anonymous):

theta^2 * sqrt(theta^2 + 9)

OpenStudy (anonymous):

very good so, arclength = ∫θ^2 sqrt(9 + θ^2) dθ, from 0 to pi, which is approximately 39.77

OpenStudy (anonymous):

Just got it. Thank you!

OpenStudy (anonymous):

as for horizontal and vertical tangent, do you know how to find them?

OpenStudy (anonymous):

Yes, I do.

OpenStudy (anonymous):

excellent.